# Sparse Diffusion Policy: A Sparse, Reusable, and Flexible Policy for Robot Learning

### v1

@@ 1 Introduction
Generalist robots are gaining substantial attention in both academia and industry, capable of performing a wide range of tasks and continually learning new ones without losing previously acquired skills [1, 2, 3, 4, 5, 6]. Traditional approaches often rely on a universal and monolithic policy [1, 2] for all tasks, activating all the parameters in the large network for even simple tasks like pushing. Besides, given the diverse nature and lifelong requirements of robot learning tasks [7, 8], when encountering a new task, these approaches typically require costly fine-tuning [9] , which carries the risk of catastrophic forgetting of previously acquired skills. For task-specific adapters such as LoRA [10], it entails expanding active parameters during inference. An alternative approach is to train separate policies for different tasks, though this requires independent and from-scratch training for each task and prevents knowledge transfer across tasks.
@@ 3 Method
Our approach integrates Mixture of Experts (MoE) layers into a transformer-based diffusion policy network [24], combined with a specifically designed training and application strategy for multitask and continual learning in robotics. Owing to the network’s structural sparsity, we refer to our method as Sparse Diffusion Policy (SDP). In the subsequent section, we first outline the problem formulation for multitask and continual imitation learning. We then discuss the integration of the Mixture of Experts (MoE) structure and explore how its sparsity, flexibility, and reusability can be specifically utilized for robot learning. Finally, we present the training strategies we have developed to further unleash their potential in the domain of robot learning.
@@ 3.3 Training Objective
where we assume that each task is equally important, i.e., . Thus, the total training objective is
@@ 4.1 Multitask Learning
In this section, we discuss sparsity and reusability of our SDP, highlighting its ability to simultaneously acquire multiple experts efficiently while sharing experts across tasks. As shown in Figure 4, we conducted three multitask learning experiments across different simulation settings: a 2D vision-based simulation, a 3D point cloud-based simulation, and real-robot experiments.
@@
2D Simulation Results. We evaluate the performance of SDP in Mimicgen [74]. Mimicgen includes 1K-10K human demonstrations per task with broad initial state distributions, effectively showing the generalization for multitask evaluation. To our knowledge, this is the first work to explore multitask training on the Mimicgen benchmark. We choose task-conditioned diffusion (TCD) [76, 19], fine-tuning the last action prediction head (TH), fine-tuning last three diffusion policy transformer layers (TT w/3 layers) as our baselines. All three baseline models require the activation of all network parameters, exhibiting a dense structure rather than a sparse one (Ours).
@@
Each model is trained for 300 epochs using one A6000 GPU for 130-150 hours and evaluated every 50 epochs. We report the average success rate of the best three checkpoints in Table 1. Our SDP surpasses all baseline models in nearly all individual tasks, achieving this with the same level of active parameters in the policy network, thanks to the sparsity of our framework. Furthermore, we have illustrated a task-expert relationship map in Figure 4. This visualization shows that each task activates only a subset of the experts within the Mixture of Experts (MoE) model, contributing to computational efficiency. More importantly, the map reveals that different tasks can activate the same expert, demonstrating the reusability of experts across various tasks.
@@
3D Simulation Results. We evaluate our approach in 3D point could-based robot learning simulation [75]. Specifically, we integrate the Mixture of Experts (MoE) into the Feed-Forward Network (FFN) blocks of the 3D diffusion policy network [75]. The number of parameters for the active experts is set to be equivalent to those in the original network. We select the toilet, faucet, and laptop tasks from the DexArt benchmark [77], with only 20-40 human demonstrations and point cloud inputs. Due to the limited number of human demonstrations, we extend the training process to 6000 epochs and evaluate the results every 200 epochs. Similar in 2D simulation, we report the average success rate of the best three checkpoints. As shown in Table 2, our method outperforms the baseline, demonstrating the effectiveness of our SDP in multitask 3D perception settings.
@@
Real Robot Experiment Results. To further evaluate the effectiveness of our SDP framework, we conduct real robot experiments on FANUC LRMate 200iD/7L robotic arm outfitted with an SMC gripper. We choose three diverse and universal tasks for evaluation: pulling a circle, picking and placing a cup, and hanging a cup. For the first two tasks, we collect 20 demonstrations each, and for the last task, we collect 40 demonstrations with two distinct hanging sticks. More details can be found in the Appendix A.2. The robot is manipulated using admittance control [78], which can achieve compliant robot motion to ensure safety during manipulation. We use TCD [76, 19] as our baseline and both models are trained for 2000 epochs. As shown in Table 3, our method greatly outperforms TCD. Additionally, we found that TCD cannot clearly distinguish the multimodal action distributions across different tasks due to the lack of sparsity, leading to model collapse. More details and visualizations can be found in the Appendix A.3.
| Method | Active Params | Square | Stack | Coffee | Hammer | Mug | Nut | Stack three | Thread | Avg. |
| TH | 52.6 M | 0.76 | 0.98 | 0.72 | 0.97 | 0.63 | 0.52 | 0.73 | 0.55 | 0.73 |
| TT w/ 3Layer | 52.6 M | 0.73 | 0.95 | 0.76 | 0.99 | 0.65 | 0.49 | 0.68 | 0.59 | 0.73 |
| TCD [76, 19] | 52.7 M | 0.63 | 0.95 | 0.77 | 0.92 | 0.53 | 0.44 | 0.62 | 0.56 | 0.68 |
| SDP(Ours) | 53.3 M | 0.74 | 0.99 | 0.83 | 0.98 | 0.70 | 0.42 | 0.76 | 0.65 | 0.76 |
@@ 4.2 Continual Learning
In this section, we evaluate the performance of the Sparse Diffusion Policy (SDP) in continual learning, specifically focusing on the flexibility of our framework. More specifically, when integrating a new task, we introduce a small, fixed number of new experts and a new router into the Mixture of Experts (MoE) structure, while freezing the existing experts, routers, and other modules (excluding the MoE). We then train only the newly-added router and experts, which enhances the efficiency of learning new tasks and avoids catastrophic forgetting [79] of previous tasks. We choose three tasks from robimimic [80] for evaluation: Can (picking and placing a can), Lift (lifting a cube), and Square (inserting a square peg into the correct post). Initially, we train our model on Can (Stage 1), then sequentially transfer to Lift (Stage 2) and Square (Stage 3). At each stage, we evaluate the performance of the current task as well as the previously learned tasks.
Comparison with baselines. First, we compare our approach with LoRA [10] and a fully fine-tuned (FFT) version of our method. Each method is trained for 500 epochs and evaluated every 50 epochs. We report the average success rate of the best three checkpoints. The results, presented in Table 4, demonstrate that our method generally exhibited superior performance in both new tasks and previous tasks. Conversely, the FFT method showed significant forgetting of previously learned tasks upon learning new ones, while the LoRA method struggled with improving the performance in the new task, particularly evident in the Square task. Moreover, the active parameters of our model remain constant during continual learning, which is significantly fewer than those of LoRA as tasks are added. These findings underscore the flexibility of our method in supporting efficient and effective continual learning.
Ablation study on continual learning. We further conducted ablation studies in the context of continual learning, presenting our results on the Square task (Stage 3). This task was chosen because it is not only the most challenging but also the final task, making it difficult to transfer knowledge to new, unseen tasks without the potential negative influence of past experiences. The results, presented in Figure 5, highlight the impact of various components of our method on the success rate. The baseline model (Ours) achieves the highest success rate. The Ours-MI_Loss variant, which excludes the Mutual Information loss, shows a drop in performance. We argue that the router tends to prefer the old experts since newly-added experts that have not yet learned anything will output random actions. And it will enforce the preference to the old frozen experts. However, the Mutual Information loss compels the router to select task-specific experts, thereby avoiding the negative effects associated with the old experts. These results underscore the importance of using a comprehensive set of MoEs and the Mutual Information loss for achieving optimal performance in complex continual learning scenarios.
| Policy Decoder | Stage 1 | Stage 2 | Stage 3 | ||||||
@@
| LoRA | 0.94 | 9.0 M | 0.94 | 1.00 | 12.0 M | 0.94 | 1.00 | 0.73 | 14.9 M |
| MOE (Ours) | 0.96 | 9.2 M | 0.94 | 1.00 | 9.2 M | 0.94 | 1.00 | 0.75 | 9.2 M |
4.3 Task Transfer
In this section, we test the reusability of the SDP, highlighting its ability to transfer to new tasks while retaining accumulated old skills. We first set up a novel experiment that rapidly transfers to a new task using only the old tasks. Then, we visualize the horizon-expert map to observe the role of old skills in transferring to new skills.
Task transfer novel experiments. To test the reusability of our model, we include a new task for transfer and use base tasks to learn the base policy. We first select the Coffee and Mug Cleanup tasks from the MimicGen dataset [74] as our base tasks. For the new task, we choose Coffee Preparation from the MimicGen dataset, which is a long-horizon and partially unseen task requiring skills from the base tasks. Notably, Coffee Preparation includes picking coffee from the drawer, whereas the base tasks involve picking coffee from the table and placing it in the drawer. This indicates that Coffee Preparation requires skills from the base tasks but not the exact same actions.
| Method | Trainable Params | Coffee Preparation |
@@ 4.3 Task Transfer
We pretrain our model on base tasks, and for the new policy, we only fine-tune the router in MoE and unfreeze the vision encoder for only 100 epochs on the Coffee Preparation task. We compare our model to a model trained directly on the Coffee Preparation task from scratch for 100 epochs. The results are shown in Table 5. They show that for the policy network, compared with the train-from-scratch method, we use less than 0.5% of the parameters and achieve better results, indicating that the reused skills from the base two tasks are significantly useful.
Skills Visualisation. As shown in Figure 6, we visualize the horizon-expert map, which indicates the activated experts at different horizon timesteps. After only tuning the router, we find that the most frequently activated experts have already been learned in the base tasks. Therefore, this verifies the reusability of the SDP model, which can cover the old skills through the learned experts. Therefore, in this setting, the experts become a pool of skills, and the router function becomes a chain planner, then SDP can acquire new skills rapidly.
@@ 5 Discussion and Limitation
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
@@ Acknowledgments
Appendix A Experiment Details
A.1 Implementation Details
For multitask learning in 2D simulation environments, we train the policy for 300 epochs with 12 layers of transformer blocks and 512 embedding dimensions. The training batch size is set to 64, and the learning rate is 0.0001, using the Adam optimizer [81]. For the evaluation, the observation steps are set to 2, planning involves 8 action steps, and we execute the first step. For SDP, we set the number of experts to 8 and activate 2 experts for each single task.
@@ A.1 Implementation Details
A.2 Real World Robot Experiments Extension
In this section, we extend the details for the real-world robot experiments, as shown in Figure 7. For the pull task, we collect 20 human demonstrations where the objective is to pull the circle to the red region in the center of the table. For the pick-and-place task, we also collect 20 demonstrations, where the task is to pick up the cup and place it on the plate. For the last task, the hang task, we collect 20 demonstrations for goal point 1 and 20 demonstrations for goal point 2. The model distinguishes between the two goal points using a learnable Fourier Embedding. After collecting all the demonstrations, the model undergoes multitask training for the three tasks simultaneously. The results are used to test which model is better for multitask learning, as shown in Table 3.
A.3 Model Collapse
In this section, we will introduce an interesting observation from our real-world robot experiments, known as ’model collapse,’ which is closely related to sparsity. As shown in Figure 8 , the TCD method completely fails in the pick-and-place task. It cannot distinguish the action space among different tasks, resulting in the execution of actions not demonstrated by humans in this task.
A.4 Continue Learning Extensive Experiments
In this section, we conduct additional experiments on continual learning by using a fully finetuned visual encoder. This approach may lead to forgetting old tasks, so we test the transfer ability of different models under these conditions. As shown in Table 6, after transferring the last task (square), our model outperforms the LoRA method in terms of transfer ability with the fully finetuned visual encoder.

### v2

@@ 1 Introduction
Generalist robots are gaining substantial attention in both academia and industry, capable of performing a wide range of tasks and continually learning new ones without losing previously acquired skills [1, 2, 3, 4, 5, 6]. Traditional approaches often rely on a universal and monolithic policy [1, 2] for all tasks, activating all the parameters in the large network for even simple tasks like pushing. Besides, given the diverse nature and lifelong requirements of robot learning tasks [7, 8], when encountering a new task, these approaches typically require costly fine-tuning [9] , which carries the risk of catastrophic forgetting of previously acquired skills. Task-specific adapters, such as LoRA [10], entail expanding active parameters during inference. An alternative approach is to train separate policies for different tasks, though this requires independent and from-scratch training for each task and prevents knowledge transfer across tasks.
@@ 3 Method
Our approach integrates Mixture of Experts (MoE) layers into a transformer-based diffusion policy network [24], combined with specifically designed training and application strategies for multitask and continual learning in robotics. Owing to the network’s structural sparsity, we refer to our method as Sparse Diffusion Policy (SDP). In the subsequent sections, we first outline the problem formulation for multitask and continual imitation learning. We then discuss the integration of the Mixture of Experts (MoE) structure and explore how its sparsity, flexibility, and reusability can be specifically utilized for robot learning. Finally, we present the training strategies we have developed to further unleash the potential in the domain of robot learning.
@@ 3.3 Training Objective
where we assume that each task is equally important, i.e., . Implementation details for are presented in Appendix A. Thus, the total training objective is
@@ 4.1 Multitask Learning
In this section, we discuss sparsity and reusability of our SDP, highlighting its ability to simultaneously acquire multiple experts efficiently while sharing experts across tasks. As shown in Figure 4, we conducted three kinds of experiments: 2D vision-based simulations, 3D point cloud-based simulations, and real-robot experiments. Implementation details are presented in Appendix B.1.
| Method | TP(M) | AP(M) | Square | Stack | Coffee | Hammer | Mug | Nut | Stack Three | Thread | Avg. |
| TH | 52.6 | 52.6 | 0.76 | 0.98 | 0.72 | 0.97 | 0.63 | 0.52 | 0.73 | 0.55 | 0.73 |
| TT w/ 3Layer | 144.7 | 52.6 | 0.73 | 0.95 | 0.76 | 0.99 | 0.65 | 0.49 | 0.68 | 0.59 | 0.73 |
| TCD [76, 19] | 52.7 | 52.7 | 0.63 | 0.95 | 0.77 | 0.92 | 0.53 | 0.44 | 0.62 | 0.56 | 0.68 |
| Octo [77] | 48.4 | 48.4 | 0.68 | 0.96 | 0.72 | 0.97 | 0.48 | 0.32 | 0.72 | 0.64 | 0.69 |
| SDP | 126.9 | 53.3 | 0.74 | 0.99 | 0.83 | 0.98 | 0.70 | 0.42 | 0.76 | 0.65 | 0.76 |
| Light SDP | 53.3 | 38.7 | 0.75 | 0.96 | 0.83 | 0.97 | 0.55 | 0.50 | 0.74 | 0.73 | 0.75 |
| Method | DexArt [78] | Adroit [79] | ||||||
| Toilet | Faucet | Laptop | Avg. | Door | Hammer | Pen | Avg. | |
| TT w/ 1Layer | 0.73 | 0.35 | 0.85 | 0.64 | 0.63 | 0.92 | 0.54 | 0.70 |
| TCD [76, 19] | 0.72 | 0.33 | 0.80 | 0.62 | 0.63 | 0.83 | 0.42 | 0.63 |
| SDP | 0.75 | 0.43 | 0.82 | 0.67 | 0.70 | 0.97 | 0.58 | 0.75 |
@@
2D Simulation Results. We evaluate the performance of SDP on 8 tasks in Mimicgen [74]. Mimicgen includes 1K-10K human demonstrations per task with broad initial state distributions, effectively showing the generalization for multitask evaluation. To our knowledge, this is the first work to explore multitask training on the Mimicgen benchmark. We choose task-conditioned diffusion (TCD) [76, 19], fine-tuning the action head (TH), fine-tuning last three transformer layers (TT w/ 3Layer), Octo [77] as our baselines. TCD and Octo require the activation of all network parameters, exhibiting a dense structure rather than a sparse one (Ours). We train our standard SDP model as well as a smaller version, referred to as Light SDP.
@@
Each model is trained for 300 epochs using one A6000 GPU for 130-150 hours and evaluated every 50 epochs. We report the average success rate of the best three checkpoints in Table 1. Thanks to the sparsity, our SDP outperforms all baselines with the same level of active parameters in the policy network. Additionally, our Light SDP also surpasses all baselines with significantly fewer active parameters, while maintaining the same level of total parameters. Another observation is that sparse models (SDP, TH, TT) outperform dense models (TCD and Octo). This suggests that dense structures may hinder the learning of distinct policies across tasks, whereas sparse structures, which allocate separate parameters for different tasks, facilitate the learning of more diverse actions. Task-expert frequency map in Figure 4 shows that each task activates only a subset of the experts, contributing to computational efficiency. More importantly, the map reveals that different tasks can activate the same expert, demonstrating the reusability of experts across various tasks.
@@
3D Simulation Results. We evaluate our approach on 3 tasks in DexArt [78] and 3 tasks in Adroit [79]. Specifically, we follow [75] and integrate the Mixture of Experts into the Feed-Forward Network blocks of the 3D diffusion policy network [75]. The number of active parameters is set to be equivalent to the original. We train for 6000 epochs and evaluate every 200 epochs. We report the average success rate of the best three checkpoints. As shown in Table 2, our method outperforms the baseline, showing the effectiveness of our SDP in 3D perception settings.
@@
Real Robot Experiment Results. We conduct real robot experiments on FANUC LRMate 200iD/7L robotic arm outfitted with an SMC gripper. We choose three diverse and universal: pulling a circle, picking and placing a cup, and hanging a cup. For the first two tasks, we collect 20 demonstrations each, and for the last task, we collect 40 demonstrations with two distinct hanging sticks. More details can be found in the Appendix B.2. The robot is manipulated using admittance control [80], which can achieve compliant robot motion to ensure safety during manipulation. The number of training epoch is 2000. As shown in Figure 5, our method greatly outperforms TCD [76, 19]. Additionally, we found that TCD cannot clearly distinguish the multimodal action distributions across different tasks. We hypothesize that this may be caused by insufficient sparsity. More details and visualizations can be found in the Appendix B.2.
@@ 4.2 Continual Learning
In this section, we evaluate the performance of the Sparse Diffusion Policy (SDP) in continual learning, with a focus on the framework’s flexibility. When integrating a new task, we add a small number of experts and a new router to the Mixture of Experts (MoE) structure, while freezing the existing experts, routers, and other non-MoE modules. We then train only the newly added router and experts, enhancing the efficiency of learning new tasks and preventing catastrophic forgetting [81] of previously learned tasks. Assume we continuously learn three tasks, . We first train the model on task (Stage 1), then on task (Stage 2), and finally on task (Stage 3), evaluating the performance on both the current and previous tasks.
Comparison with baselines. We choose three tasks from robimimic [82] for evaluation: Can Lift Square. First, we compare our approach with LoRA [10] and a fully fine-tuned (FFT) version of our method. Each method is trained for 500 epochs and evaluated every 50 epochs. We report the average success rate of the best three checkpoints. The results, presented in Table 3, demonstrate that our method generally exhibited superior performance in both new tasks and previous tasks. Conversely, FFT exhibited significant forgetting of previously learned tasks when learning new ones, whereas LoRA struggled with a substantial increase in the number of active parameters. These findings underscore the flexibility of our SDP.
Ablation study on continual learning. In this section, we report ablation study on mutual information (MI) loss in Section 3.3. We selected three tasks from MimicGen [74]: Stack Hammer Cleanup Coffee, where task complexity progressively increases (e.g., the Coffee task requires not only picking and placing but closing the coffee cap.). Since our SDP avoids catastrophic forgetting [81], we report only the performance of the newly added tasks in Figure 6. The results emphasize the significance of MI loss. We argue that the router initially favors previously trained experts, as newly added experts, being untrained, produce random actions, reinforcing reliance on frozen experts. However, MI loss encourages the router to select task-specific experts, promoting the use of seldom-selected experts. Additional ablation studies are presented in Appendix C.3, where experiments further demonstrate the critical role of comprehensive MoE structures and MI loss in achieving optimal performance in complex continual learning settings.
| Method | Stage 1 | Stage 2 | Stage 3 | ||||||
@@
| LoRA[10] | 0.94 | 9.0 M | 0.94 | 1.00 | 12.0 M | 0.94 | 1.00 | 0.73 | 14.9 M |
| SDP (Ours) | 0.96 | 9.2 M | 0.94 | 1.00 | 9.2 M | 0.94 | 1.00 | 0.75 | 9.2 M |
4.3 Efficient Task Transfer
In this section, we evaluate the reusability of our SDP, highlighting its ability to efficiently transfer to new tasks by learning how to leverage previously acquired experts (router tuning). Additionally, we visualize the evolution of expert scores to assess the role of prior skills in enabling the transfer to new tasks.
Efficient task transfer experiment. We first train our SDP on base tasks, then freeze all experts and fine-tune only the router for the new task. Specifically, we use Coffee and Mug Cleanup as the base tasks and Coffee Preparation as the new task, all from MimicGen [74]. Coffee Preparation is a long-horizon, partially unseen task requiring unique skills, such as moving the mug to the drip tray. After pretraining on the base tasks, we fine-tune only the MoE router and the vision encoder for the new task. For comparison, we also train a model with the same structure from scratch. Both models are trained for 100 epochs. As shown in Table 4, our method uses less than 0.5% of the parameters compared to the train-from-scratch approach while achieving better performance, demonstrating that the experts learned from the base tasks effectively cover diverse skills and offer strong representation for distinct tasks.
| Method | Train. Params | Success Rate |
@@ 4.3 Efficient Task Transfer
Skills visualisation. As shown in Figure 7, we visualize the evolution of expert scores on the Coffee, Mug Cleanup, and Coffee Preparation tasks. We observe that when knowledge from Coffee is required but not present in Mug Cleanup, the experts trained on Coffee are activated. Four typical examples are illustrated on the right of Figure 7. By composing these experts, SDP can acquire new skills, such as moving the mug to the coffee machine’s drip tray. Additionally, we find that some experts are consistently activated in Coffee Preparation, while others are seldom used. This insight suggests the possibility of policy distillation, which we leave for future work. More experiment details and explanations are presented in Appendix D.
@@ 5 Discussion and Limitation
Appendix A Mutual Information Loss
In order to calculate , we need to get , and . can be calculated by counting the number of selection times of expert across all the tasks. We assume each task is equally important so where is the number of tasks. For , we have . can be calculated by counting the number of selection times of expert for task .
Appendix B Multitask Learning Experiments
B.1 Implementation Details
For multitask learning in 2D simulation environments, we train the policy for 300 epochs using 12 transformer blocks with 512 embedding dimensions. The batch size is set to 64, and the learning rate is 0.0001, optimized with Adam [83]. During evaluation, we use 2 observation steps, followed by 8 action planning steps, executing only the first step. For SDP, we use 8 experts, activating 2 experts per task, with each expert having the same size as the original Feed-Forward Network (FFN). In Light SDP, we increase the number of experts to 16 and activate 8 experts, with the size of each expert reduced to of the original FFN to maintain the overall model size. For Octo [77] structure, we reimplemented it with the same design, using a learnable task embedding as the transformer input token and a diffusion policy head.
@@ B.1 Implementation Details
B.2 Real Robot Experiments
In this section, we provide additional details on the real-world robot experiments, as depicted in Figure 9. For the pull task, we collected 20 human demonstrations, where the objective is to pull the circle to the red region at the center of the table. For the pick-and-place task, we also gathered 20 demonstrations, where the goal is to pick up the cup and place it on the plate. Lastly, for the hang task, we collected 20 demonstrations for each of the two goal points. The initial object position is used as the state input for the Pull and Pick-and-Place tasks, while the goal point serves as the state input for the Hang task. All state inputs are encoded using learnable Fourier embeddings.
We observe that TCD [76, 19] struggles to capture the multimodality of the task-specific policies and distinguish between different task behaviors, leading to failures. In detail, TCD always rotates the end effectors in the Pick and Place (See Figure 8), a behavior not observed in the demonstrations. Instead, rotations are necessary only for Hang. This observation indicates that TCD tends to conflate policies from different tasks and struggles to capture multimodal action distributions across diverse tasks.
Appendix C Continual Learning Experiments
C.1 Implementation Details
@@
C.2 Comparison with Baselines
In this section, we conduct additional continual learning experiments using a fully fine-tuned visual encoder. As shown in Table 5, after transferring to the final task (Square), our model outperforms the LoRA method in terms of transferability with the fully fine-tuned visual encoder. However, fully fine-tuning the visual encoder results in complete forgetting of previous tasks. Therefore, utilizing a comprehensive set of MoEs is essential for effective continual learning.
C.3 Ablation Study
In this section, we present ablation study results on three easy tasks: Can Lift Square. Can involves picking and placing a can, Lift focuses on lifting a cube, and Square requires inserting a square peg into the correct post. We report the performance on the Square task at Stage 3 in Figure 10. These results highlight the importance of employing a comprehensive set of MoEs and the Mutual Information loss to achieve optimal performance in complex continual learning scenarios. Compared to the results in Figure 6, we observe that in more challenging continual learning settings (Stack Hammer Cleanup Coffee), Mutual Information loss plays a more significant role in learning new tasks.